# Single-Shot Readout and Weak Measurement of a Tin-Vacancy Qubit in Diamond

### v1

@@ II.3 Spin polarization fidelity
Measurement of spin polarization, as in Fig. 2d, also bounds the polarization fidelity. This quantifies the extent to which a resonant drive prepares the qubit in the pure states $\ket{\downarrow}$ or $\ket{\uparrow}$, and is defined as $F_{\mathrm{pol}}=1-\frac{1}{2}\Gamma_{\mathrm{min}}/\Gamma_{\mathrm{max}}$, where $\Gamma_{\mathrm{max}}$ is the maximum count rate and $\Gamma_{\mathrm{min}}$ is the mean count rate in the $t\gg 0$ limit [22, 25]. We measure $F_{\mathrm{pol}}=92.8\%$ at $\zeta=147\degree$, and $F_{\mathrm{pol}}=94.9\%$ at $\zeta=53\degree$, consistent with other examples of high-fidelity spin polarization of ${}^{-}$’s [22, 25, 39, 26, 27]. This is limited by non-zero $\Gamma_{\mathrm{min}}$, which is dominated by scattering from the excitation laser into the collection path. This technical source of noise can be mitigated by better spectral filtering, and thus our polarization fidelity measurement is a lower bound.
@@ II.4 Spin control
We now characterize performance of the ${}^{-}$ as a spin qubit. When operated at maximum cyclicity ($\zeta=147\degree$) we demonstrate high-fidelity spin manipulation using a microwave control pulse, Fig. 2f. At this operating point we demonstrate a $\pi$-pulse time of 80 ns, operated at $\omega_{q}/2\pi=3.677\,\mathrm{GHz}$ (with $|\vec{B}|=125\,\mathrm{mT}$) and using microwave input power and packaging similar to Ref. [26].
In Fig. 2g we also measure the Rabi rate (defined as the inverse of twice the $\pi$-pulse time), as a function of $\zeta$. To avoid frequency dependent changes to microwave power delivery, we keep the qubit frequency fixed at $\omega_{q}/2\pi=4.000\pm 0.002\,\mathrm{GHz}$ by changing the amplitude $|\vec{B}|$ at each value of $\zeta$. Data is compared to a model (from Ref. [26]) that assumes a drive of $|\vec{B}|=0.6\,\mathrm{mT}$ at the spin location, and which is oriented perpendicular to $\vec{\mu}$. For this drive orientation, the Rabi rate is expected to be highest near $\approx 147\degree$ where $\vec{B}$ and $\vec{\mu}$ are maximally aligned, and symmetric with angle around this operating point. Disagreement between theory and experiment may be due to drifting microwave power delivery, and/or, miscalibration of one or both magnetic coils. However, the main result of Fig. 2g is that microwave Rabi rate changes little with $\zeta$. Therefore, at least for this level of strain, ${}^{-}$ experiments can expect to operate at optimal cyclicity without compromise to microwave spin control.
In summary, we have determined that resonant excitation polarizes the ${}^{-}$ spin state while inducing photon emission. These effects have a strong dependence on the alignment between the spin dipole and external magnetic field, with closer alignment leading to slower polarization, higher cyclicity, and greater emission. We also demonstrate that high cyclicity is compatible with rapid microwave spin control. For the remainder of this work we operate at $\zeta=147\degree$ where cyclicity is maximized for this experimental setup.
@@ III Readout characterization
Figure 3: Single-shot readout of a single electron spin. (a) Readout is characterized using two waveforms (wfm-A and wfm-B) which are each repeated for $4.27\times 10^{5}$ cycles. Each cycle begins with preparation of the spin in the polarized states $\ket{\uparrow}$ or $\ket{\downarrow}$ (wfm-A or wfm-B, respectively). The spin state is then projectively measured twice. Between measurements the spin is either coherently flipped by a microwave $\pi$-pulse (wfm-A), or, no pulse is applied (wfm-B, where readout of spin-down state already flips the spin). Each cycle also includes two charge resonance checks (CRCs; not shown, see Fig. 11 for details). (b,c) Count distributions of the first and second readouts, respectively. Lower (higher) number distributions have a mean count number of $\bar{n}_{d}$ ($\bar{n}_{b}$) and are associated with the projection of the qubit in the $\ket{\uparrow}$ ($\ket{\downarrow}$) states prior to measurement. A given readout event yields the result “1” (“0”) if the counts collected during this event fall below (at or above) a threshold $N_{r}=1$, illustrated here by the black dashed lines. Here, both readout steps have a fidelity of $F_{r}=87.4\%$ (Eq. 3) and (d) a conditional fidelity of $F_{c}\approx 98.5\%$ (Eq. 4). Note that data in (b,c,d) has been post-selected for cycles which pass both CRCs (7.7% of cycles for wfm-A, and 7.4% of cycles for wfm-B, see Fig. 12 for details).
@@ III.1 Readout fidelity
Given our understanding of spin polarization and cyclicity, we now characterize readout of the ${}^{-}$ spin state. To do so, we measure the time-tagged photon statistics of two subsequent readout steps, Fig. 3a. Depending on spin state preparation, the histogram of counts collected over many readout events in a set window follows either a “bright” distribution with a mean photon number of $\bar{n}_{b}>0$, or a “dark” distribution with a mean photon number $\bar{n}_{d}\approx 0$.
Readout fidelity is defined as [42]:
$F_{r}=1-\frac{1}{2}P(1|\downarrow)-\frac{1}{2}P(0|\uparrow),$ (3)
@@
Readout results are shown in Fig. 3b,c. After spin state initialization, two subsequent readout steps separated by a microwave $\pi$-pulse (“wfm-A”), or with no pulse (“wfm-B”). In each readout step the resonant laser is pulsed for $100\,\mathrm{\mu s}$, long enough to fully polarize the spin state. However emitted counts are only recorded for the first $50\,\mathrm{\mu s}$ of this drive, since after this time the noise (predominantly scatter from the excitation laser) can exceed the emitted signal and thus reduce fidelity. For both the first and second readout steps we measure a readout fidelity of $F_{r}=87.4\%$. This is in the “single-shot” regime, meaning a single projective measurement, or experimental “shot”, correctly determines the qubit with this fidelity. For this measurement, the “dark” distribution has $\bar{n}_{d}\leq 0.2$ counts per cycle, and “bright” distributions have $\bar{n}_{b}\approx 4$ counts per cycle.
Both before and after these readout steps, we include charge resonance checks [42] (CRCs, not shown in Fig. 3a). Each CRC is a spin-independent check that consists of simultaneously driving both spin-preserving transitions (A1 and B2), and recording the check as a pass only if greater than a certain number of counts are collected during the check (see Appendix C.3.2 for details). Importantly, CRCs allow us to selectively characterize our qubit only when it is in the correct charge state and not suffering from spectral diffusion. For example cycles are not recorded when the qubit “blinks off” – an erasure error that can be caused by the excitation laser changing the defect’s charge state. By setting a CRC pass threshold to maximize readout fidelity while keeping a significant fraction of cycles, we retain $7.4\%$ ($7.7\%$) of cycles of wfm-A (wfm-B), out of $4.27\times 10^{5}$ cycles total for each waveform. The data shown in Fig. 3 includes passing cycles, only. See Fig. 13 for analysis of readout as a function of pass threshold.
@@ III.2 Conditional fidelity
Readout can be further characterized by a “conditional fidelity”, $F_{c}$, which conditions on the results of two subsequent measurements. For example, consider the characterization sequence described in Fig. 3a. The result of these two measurements should be anti-correlated due to the insertion of the spin $\pi$-pulse during the second readout, enabling the extraction of the conditional fidelity. We therefore expect results “1” then “0” for wfm-A, and results “0” then “1” for wfm-B. Using the data in Fig. 3d we find a conditional fidelity of $F_{c}=98.5\%$, defined as [43, 44]:
@@
Here, $i,j\in\{0,1\}$ are measurement results with $i\neq j$, and $P(j|i)\geq P(i|j)$ are the conditional probabilities of result $i$ ($j$) in the second readout given result $j$ ($i$) in the first readout. Like CRCs, conditional fidelity provides an avenue to retain data to only include select experimental cycles. This boosts the fidelity of selected measurements at the cost of a reduced rate of data acquisition.
III.3 Quantum non-demolition fidelity
Correlating the results of two subsequent measurements also quantifies the degree to which readout is quantum non-demolition (QND) [45, 46]. In quantum theory, a qubit which has been projectively measured should remain in its measured eigenstate [47]. This is a desirable property; for example, it is needed in quantum error correction algorithms that require an ancilla qubit to be repeatedly measured.
However, a realistic quantum measurement is not always QND. For example, a measurement that correctly determines a color center’s spin state, but in the process kicks the color center into a different charge state and thus out of the qubit subspace (erasure error), is not QND. A measurement which heats the qubit, causing bit flip errors independent of the measurement result is also not QND.
The ${}^{-}$ qubit and many other atom-like systems have the property that, due to spin polarization under resonant drive, when the qubit is measured in the spin state associated with bright emission its eigenstate is also flipped. Therefore the ${}^{-}$ readout described in this work is not QND by the definition $F_{\mathrm{QND}}=\left[P(1|1)+P(0|0)\right]/2$ [45], where $P(1|1)$ and $P(0|0)$ are the conditional probabilities that subsequent measurements each yield the same result. However, we may define an alternate definition that captures the extent to which a qubit remains in the eigenstate it should be in after measurement (i.e. is not subject to erasure error, or an unexpected bit-flip error). We define this QND-equivalent fidelity, $F_{q}$, as:
@@ III.3 Quantum non-demolition fidelity
where $i,j\in\{0,1\}$ and $i\neq j$. Here, $P(i|j)$ and $P(j|i)$ are the conditional probabilities that subsequent readout results are anti-correlated as they are expected to be for the experimental sequence in Fig. 3a. Eq. 5 therefore captures the essence of QND fidelity; the extent to which readout does not unexpectedly perturb the qubit. After all, a measurement result which deterministically flips the qubit state can in-principle be corrected by a $\pi$-pulse to return the qubit to its measured state, and would thus be QND by the standard definition [45].
Using Eq. 5, we compute $F_{q}=76.3\%$ for the data in wfm-A and $F_{q}=77.9\%$ for the data in wfm-B. Post-measurement, the qubit therefore remains in its expected eigenstate a considerable fraction of the time. Infidelity is dominated by cycles in which both readouts yield a dark result, which may result from blinking of the qubit during readout even for cycles when both CRCs are passed, or, from the non-zero overlap of the bright and dark distributions. Finding ways to increase $F_{q}$ is an important research direction for future experiments and is key to implement many quantum error correction protocols.
Figure 4: Measurement dependence on power and efficiency. (a) Readout performance is characterized as a function of measurement power $p$. Data is post-selected on a charge resonance check (CRC) before each cycle, not shown. (b) Mean counts $\bar{n}_{b}$ and $\bar{n}_{d}$ in the bright or dark distributions, respectively, plotted vs. normalized power $p/p_{\mathrm{sat}}$, where $p$ is the optical drive power and $p_{\mathrm{sat}}$ is the on-resonance saturation power. The difference $\bar{n}_{b}-\bar{n}_{d}$ captures signal from emitted photons, and is fit to Eq. 6 to obtain a saturation power of $p_{\mathrm{sat}}=313\pm 8\,\mathrm{nW}$ and a measurement efficiency of $\eta=0.1\%$. (c) Simulated readout infidelity $1-F_{r}$, plotted on a log scale. Here $F_{r}$ is simulated using Eq. 28 which assumes emission to follow Poissonian statistics, and assumes $\chi=2.2\times 10^{3}$ as in experiment. Greater efficiency lowers the drive power needed to achieve high fidelity.
@@ III.4 Efficiency
The electron spin readout described in Fig. 3 is not perfect, as evidenced by the overlap of the distributions in Fig. 3b,c. In this section we study what limits the number of collected photons, including characterizing the “measurement efficiency”, $\eta$, that quantifies overall loss between the qubit and detector.
Here we define measurement efficiency to be equivalent to a beamsplitter with transmission $\eta$ placed between a source of emission and an ideal photon counter, Fig. 1. This is the probability of detecting a “click” per emission event, and is the product of the probability of all sources of loss. It is crucial to characterize and understand $\eta$ in order to improve future experiments by reducing loss. Lower loss (increased $\eta$) will allow for higher fidelity single-shot readout and in future quantum networks will allow improved rates of entanglement generation.
Cyclicity describes the number of times, on average, that the emitter goes through excitation / relaxation before polarizing. From Fig. 2d,e we expect $\chi\approx 2.2\times 10^{3}$ emission events per readout, assuming the duration of readout is long compared to the polarization time. We measure far fewer counts per readout because most emission events are not recorded by our detector. From the qualitative argument that $\eta\approx\bar{n}_{b}/\chi$ and $\bar{n}_{b}\approx 4$, we estimate $\eta\approx 0.2\%$ for the data in Fig. 3.
@@
which assumes preparation in the spin state associated with bright emission, and at high power simplifies to $\bar{n}_{b}-\bar{n}_{d}\approx\chi\eta$. Here, $\kappa$ is the power-dependent polarization rate, $\chi=2244\pm 108$ is the cyclicity, $\tau=50\,\mathrm{\mu s}$ is the duration of the readout collection window, and as before $\bar{n}_{b}$ ($\bar{n}_{d}$) are the mean number of counts detected when the qubit is prepared in the spin state associated with bright (dark) emission.
@@
We fit Eq. 6 to the measurement in Fig. 4b. From this fit we determine efficiency of $\eta=0.992\times 10^{-3}\pm 5.73\times 10^{-6}\approx 0.1\%$; this is lower than we expect from the data in Fig. 3 because of the lower mean count number difference of $\bar{n}_{b}-\bar{n}_{d}\approx 2$ collected photons per cycle at high powers. We attribute this difference to changing conditions of the setup between measurements including drift in alignment and fluctuating polarization of the excitation laser.
@@
Fig. 4b shows that the mean number of emitted photons plateaus near a saturation power of $p_{\mathrm{sat}}=313\pm 8\,\mathrm{nW}$, specified at the input to the cryostat. This affirms the straightforward conclusion that readout should be optimized by operating near saturation. Below saturation, emission can increased by pumping harder. Above saturation, noise may increase with power but signal will not. See Fig. 10 for a measurement of readout fidelity vs. power.
@@
Many sources of loss contribute to inefficiency including non-radiative decay, filtering of emission to collect the phonon sideband only, loss between the qubit and detector, and detector inefficiency (see Table 2 for details). However, we estimate the dominant source of loss is the scattering of emission into bulk diamond rather than into the collection path. Although this is the same emitter used in Ref. [26], from fine-tuned optimization using nano-positioners we confirm that this emitter is actually in a mesa structure (see Fig. 9) rather than a photonic nanopillar. Numerical simulations of an ${}^{-}$ center within this structure predict that only a modest $\approx 5\%$ of emitted light is routed to collection (compared to $\approx 3\%$ for an emitter in bulk).
@@
This fraction can be greatly increased by use of nanophotonics. For example, at visible wavelengths, an emitter integrated with a waveguide can allow for $\approx 45\%$ of channelling of photons into the waveguide mode [48]. For 1D photonic crystal nanobeams, this number can be increased to $\approx 90\%$ [29]. Likewise, inverse-designed grating couplers can extract $\approx 25\%$ [49] and tapered fiber coupling up to $\approx 90\%$ [50] of photons in a waveguide mode. Since stable, narrow linewidth ${}^{-}$’s have already been successfully incorporated into nanophotonic structures [28, 29, 30, 31, 32, 34, 33], utilizing nanophotonics offers a clear path toward an order-of-magnitude higher efficiency. This will allow for much higher spin readout fidelity, Fig. 4c, over expanded magnetic field orientations and strain regimes. For instance, this will enable single-shot readout of highly strained ${}^{-}$’s (with lower cyclicity), for which microwave spin control requires less drive power [27, 24].
@@
Figure 5: Control of qubit dephasing via weak quantum measurement. (a) A qubit prepared in a superposition state has an initial coherence, specified by the amplitude of its off-diagonal density matrix elements $|\hat{\rho}_{\downarrow\uparrow}|$. This is illustrated on the Bloch sphere for the superposition state (|+|$(\ket{\downarrow}+\ket{\uparrow})/\sqrt{2}$. (b) Measurement causes decoherence at rate $\Gamma_{\phi}$. In a projective measurement the qubit is entirely dephased, but more generally after a weak measurement its coherence is reduced to $|\hat{\rho}_{\downarrow\uparrow}^{\prime}|\leq|\hat{\rho}_{\downarrow\uparrow}|$. (c) This interplay is studied by inserting a weak measurement of variable power $p/p_{\mathrm{sat}}$ and variable duration $\tau$ into the center of a dynamical decoupling experiment. (See Fig. 11 for details.) The (projective) readout signal of this experiment measures the ratio $r$ of the remaining coherence to the initial coherence, $r=|\hat{\rho}_{\downarrow\uparrow}^{\prime}|/|\hat{\rho}_{\downarrow\uparrow}|$, which decays exponentially with weak measurement time as $r=\mathrm{exp}\left(-\Gamma_{\phi}\tau\right)$. (d) Selected measurements of $r$ as a function of $\tau$ and $p/p_{\mathrm{sat}}$, and compared to exponential fits. (e) Dephasing rate vs. power. Below saturation, the dephasing rate rises linearly with power. Above saturation, it plateaus to $\gamma/4$, the maximum dephasing rate set by the ${}^{-}$’s optical decay rate $\gamma$. Data are fit to Eq. 7.

### v2

@@ II.3 Spin polarization fidelity
Measurement of spin polarization, as in Fig. 2d, also bounds the polarization fidelity. This quantifies the extent to which a resonant drive prepares the qubit in the pure states $\ket{\downarrow}$ or $\ket{\uparrow}$, and is defined as $F_{\mathrm{pol}}=1-b/(2a)$, where $a$ is the maximum count rate and $b$ is the mean count rate in the $t\gg 0$ limit [26, 29]. We measure $F_{\mathrm{pol}}=92.8\%$ at $\zeta=147\degree$ and $F_{\mathrm{pol}}=94.9\%$ at $\zeta=53\degree$, consistent with other examples of high-fidelity spin polarization of SnV- centers [26, 29, 43, 30, 31]. This is limited by nonzero $b$, which is dominated by scattering from the excitation laser into the collection path.
This technical source of noise can be mitigated by better spectral filtering, and, thus, our polarization fidelity measurement is only a lower bound. With better filtering, we would expect to bound polarization fidelity at $F_{\mathrm{pol}}\gtrsim 99\%$, consistent with other SnV- results [26, 29, 43]. Polarization fidelity is eventually limited by the spectral overlap of the A1 and B2 transitions, which here are detuned by 500 MHz, an order of magnitude greater than their linewidths. It is also limited by the finite spin $T_{1}$ time, which at 1.7 K is order seconds [30, 31], much greater than the order 10 $\,\mathrm{\mu s}$ polarization time.
@@ II.4 Spin control
We now characterize performance of the SnV- as a spin qubit. When operated at maximum cyclicity ($\zeta=147\degree$) we demonstrate high-fidelity spin manipulation using a microwave control pulse, Fig. 2f. At this operating point we demonstrate a $\pi$-pulse time of 80 ns, operated at $\omega_{q}/2\pi=3.677\,\mathrm{GHz}$ (with $|\vec{B}|=125\,\mathrm{mT}$) and using microwave input power and packaging similar to Ref. [30].
In Fig. 2g we also measure the Rabi frequency (defined as the inverse of twice the $\pi$-pulse time), as a function of $\zeta$. To avoid frequency-dependent changes to microwave power delivery, we keep the qubit frequency fixed at $\omega_{q}/2\pi=4.000\pm 0.002\,\mathrm{GHz}$ by changing the amplitude $|\vec{B}|$ at each value of $\zeta$. Data are compared to a model (from Ref. [30]) that assumes a drive of $|\vec{B}|=0.6\,\mathrm{mT}$ at the spin location, and which is oriented perpendicular to $\vec{\mu}$. For this drive orientation, the Rabi frequency is expected to be highest near approximately $147\degree$ where $\vec{B}$ and $\vec{\mu}$ are maximally aligned, and symmetric with angle around this operating point. Disagreement between theory and experiment may be due to drifting microwave power delivery and/or miscalibration of one or both magnetic coils. However, the main result of Fig. 2g is that microwave Rabi frequency changes little with $\zeta$. This result is consistent with numerical simulations of the SnV- Hamiltonian (Fig. 9) and shows that for this system, high cylicity and optimal readout are not to the detriment of spin control.
In summary, we have determined that resonant excitation polarizes the SnV- spin state while inducing photon emission. These effects have a strong dependence on the alignment between the spin dipole and external magnetic field, with closer alignment leading to slower polarization, higher cyclicity, and greater emission. We also demonstrate that high cyclicity is compatible with rapid microwave spin control. For the remainder of this work, we operate at $\zeta=147\degree$, where cyclicity is maximized for this experimental setup.
@@ III Readout characterization
Figure 3: Single-shot readout of a single electron spin. (a) Readout is characterized using two waveforms (wfm-A and wfm-B) which are each repeated for $4.27\times 10^{5}$ cycles. Each cycle begins with preparation of the spin in the polarized states $\ket{\uparrow}$ or $\ket{\downarrow}$ (wfm-A or wfm-B, respectively). The spin state is then projectively measured twice. Between measurements, the spin is either coherently flipped by a microwave $\pi$-pulse (wfm-A), or, no pulse is applied (wfm-B, where readout of the spin-down state already flips the spin). Each cycle also includes two charge resonance checks (CRCs; not shown — see Fig. 12 for details). (b),(c) Count distributions of the first and second readouts, respectively. Lower (higher) number distributions have a mean count number of $\bar{n}_{d}$ ($\bar{n}_{b}$) and are associated with the projection of the qubit in the $\ket{\uparrow}$ ($\ket{\downarrow}$) states prior to measurement. A given readout event yields the result “1” (“0”) if the counts collected during this event fall below (at or above) a threshold $N_{r}=1$, illustrated here by the black dashed lines. Here, both readout steps have a fidelity of $F_{r}=87.4\%$ (Eq. 3) and (d) a conditional fidelity of $F_{c}\approx 98.5\%$ (Eq. 4). Note that data in (b)-(d) are post-selected for cycles which pass both CRCs (7.7% of cycles for wfm-A, and 7.4% of cycles for wfm-B, see Fig. 13 for details).
@@ III.1 Readout fidelity
Given our understanding of spin polarization and cyclicity, we now characterize readout of the SnV- spin state. To do so, we measure the time-tagged photon statistics of two subsequent readout steps, Fig. 3a. Depending on spin state preparation, the histogram of counts collected over many readout events in a set window follows either a “bright” distribution with a mean photon number of $\bar{n}_{b}>0$ or a “dark” distribution with a mean photon number $\bar{n}_{d}\approx 0$.
Readout fidelity is defined as [46]
$F_{r}=1-\frac{1}{2}P(1|\downarrow)-\frac{1}{2}P(0|\uparrow).$ (3)
@@
Readout results are shown in Figs. 3b and 3c. After spin state initialization, two subsequent readout steps separated by a microwave $\pi$ pulse (“wfm-A”), or with no pulse (“wfm-B”). In each readout step, the resonant laser is pulsed for $100\,\mathrm{\mu s}$, long enough to fully polarize the spin state. However emitted counts are recorded only for the first $50\,\mathrm{\mu s}$ of this drive, since after this time the noise (predominantly scatter from the excitation laser) can exceed the emitted signal and, thus, reduce fidelity. For both the first and second readout steps we measure a readout fidelity of $F_{r}=87.4\%$. This is in the “single-shot” regime, meaning a single projective measurement, or experimental “shot,” correctly determines the qubit state with this fidelity. For this measurement, dark distribution have $\bar{n}_{d}\leq 0.2$ counts per cycle, and bright distributions have $\bar{n}_{b}\approx 4$ counts per cycle.
Both before and after these readout steps, we include charge resonance checks [46] (CRCs, not shown in Fig. 3a). Each CRC is a spin-independent check that consists of simultaneously driving both spin-preserving transitions (A1 and B2), and recording the check as a pass only if greater than a certain number of counts are collected during the check (see Appendix C.3.1 for details). Importantly, CRCs allow us to selectively characterize our qubit only when it is in the correct charge state and not suffering from spectral diffusion, and, thus, Eq. 2 may be applied in the limit where drive detuning is small compared to a linewidth. Cycles are also not recorded when the qubit “blinks off”— an erasure error that can be caused by the excitation laser changing the defect’s charge state. By setting a CRC pass threshold to maximize readout fidelity while keeping a significant fraction of cycles, we retain $7.4\%$ ($7.7\%$) of cycles of wfm-A (wfm-B), out of $4.27\times 10^{5}$ cycles total for each waveform. The data shown in Fig. 3 include passing cycles, only. See Fig. 14 for analysis of readout as a function of pass threshold.
@@ III.2 Conditional fidelity
Readout can be further characterized by a “conditional fidelity” $F_{c}$, which conditions on the results of two subsequent measurements. For example, consider the characterization sequence described in Fig. 3a. The result of these two measurements should be anticorrelated due to the insertion of the spin $\pi$ pulse during the second readout, enabling the extraction of the conditional fidelity. We, therefore, expect results “1” then “0” for wfm-A, and results “0” then “1” for wfm-B. Using the data in Fig. 3d we find a conditional fidelity of $F_{c}=98.5\%$, defined as [47, 48, 49]
@@
Here, $i,j\in\{0,1\}$ are measurement results with $i\neq j$, and $P(j|i)\geq P(i|j)$ are the conditional probabilities of result $i$ ($j$) in the second readout given result $j$ ($i$) in the first readout. Like CRCs, conditional fidelity provides an avenue to retain data to include only select experimental cycles. This boosts the fidelity of selected measurements at the cost of a reduced rate of data acquisition.
III.3 Quantum nondemolition fidelity
Correlating the results of two subsequent measurements also quantifies the degree to which readout is quantum nondemolition (QND) [50, 51]. In quantum theory, a qubit which has been projectively measured should remain in its measured eigenstate [52]. This is a desirable property; for example, it is needed in quantum error correction algorithms that require an ancilla qubit to be repeatedly measured and real-time control logic to be implemented based on this measurement result [53]. Repetitive measurement may also be used to enhance signal to noise of a measurement, including via use of a nuclear spin ancilla [54].
However, a realistic quantum measurement is not always QND. For example, a measurement that correctly determines a color center’s spin state, but in the process kicks the color center into a different charge state and, thus, out of the qubit subspace (erasure error), is not QND. A measurement which heats the qubit, causing bit-flip errors independent of the measurement result, is also not QND.
The SnV- qubit and many other atomlike systems have the property that, due to spin polarization under resonant drive, when the qubit is measured in the spin state associated with bright emission its eigenstate is also flipped. Therefore, the SnV- readout described in this work is not QND by the definition $F_{\mathrm{QND}}=\left[P(1|1)+P(0|0)\right]/2$ [50], where $P(1|1)$ and $P(0|0)$ are the conditional probabilities that subsequent measurements each yield the same result. However, we may define an alternate definition that captures the extent to which a qubit remains in the eigenstate it should be in after measurement (i.e., is not subject to erasure error, or an unexpected bit-flip error). We define this QND-equivalent fidelity $F_{q}$ as
@@ III.3 Quantum nondemolition fidelity
where $i,j\in\{0,1\}$ and $i\neq j$. Here, $P(i|j)$ and $P(j|i)$ are the conditional probabilities that subsequent readout results are anticorrelated as they are expected to be for the experimental sequence in Fig. 3a. Equation 5, therefore, captures the essence of QND fidelity: the extent to which readout does not unexpectedly perturb the qubit. After all, a measurement result which deterministically flips the qubit state can, in principle, be corrected by a $\pi$ pulse to return the qubit to its measured state, and would, thus, be QND by the standard definition [50].
Using Eq. 5, we compute $F_{q}=76.3\%$ for the data in wfm-A and $F_{q}=77.9\%$ for the data in wfm-B. Postmeasurement, the qubit therefore remains in its expected eigenstate a considerable fraction of the time. Infidelity is dominated by cycles in which both readouts yield a dark result, which may result from blinking of the qubit during readout even for cycles when both CRCs are passed or from the nonzero overlap of the bright and dark distributions.
Finding ways to increase $F_{q}$ is an important research direction for future experiments and is key to implement many quantum error correction protocols. To this end, SiV- experiments that make use of cavity quantum electrodynamics (QED) to show high-fidelity QND readout [19, 20, 21, 22, 23] are a promising blueprint for future SnV- experiments.
Figure 4: Measurement dependence on power and efficiency. (a) Readout performance is characterized as a function of measurement power $p$. Data are postselected on a CRC before each cycle, not shown. (b) Mean counts $\bar{n}_{b}$ and $\bar{n}_{d}$ in the bright or dark distributions, respectively, plotted vs normalized power $p/p_{\mathrm{sat}}$, where $p$ is the optical drive power and $p_{\mathrm{sat}}$ is the on-resonance saturation power. The difference $\bar{n}_{b}-\bar{n}_{d}$ captures signal from emitted photons and is fit to Eq. 6 to obtain a saturation power of $p_{\mathrm{sat}}=313\pm 8\,\mathrm{nW}$ and a measurement efficiency of $\eta=0.1\%$. (c) Simulated readout infidelity $1-F_{r}$, plotted on a log scale. Here, $F_{r}$ is simulated using Eq. 28, which assumes emission to follow Poissonian statistics and assumes $\Lambda=2.2\times 10^{3}$ as in experiment. Greater efficiency lowers the drive power needed to achieve high fidelity.
@@ III.4 Efficiency
The electron spin readout described in Fig. 3 is not perfect, as evidenced by the overlap of the distributions in Figs. 3b and 3c. In this section, we study what limits the number of collected photons, including characterizing the “measurement efficiency” $\eta$ that quantifies overall loss between the qubit and detector.
Here we define measurement efficiency to be equivalent to a beam splitter with transmission $\eta$ placed between a source of emission and an ideal photon counter, Fig. 1. This is the probability of detecting a “click” per emission event and is the product of the probability of all sources of loss. It is crucial to characterize and understand $\eta$ in order to improve future experiments by reducing loss. Lower loss (increased $\eta$) will allow for higher fidelity single-shot readout and in future quantum networks will allow improved rates of entanglement generation.
Cyclicity describes the number of times, on average, that the emitter goes through excitation and relaxation before polarizing. From Figs. 2d and 2e we expect $\Lambda\approx 2.2\times 10^{3}$ emission events per readout, assuming the duration of readout is long compared to the polarization time. We measure far fewer counts per readout because most emission events are not recorded by our detector. From the qualitative argument that $\eta\approx\bar{n}_{b}/\Lambda$ and $\bar{n}_{b}\approx 4$, we estimate $\eta\approx 0.2\%$ for the data in Fig. 3.
@@
which assumes preparation in the spin state associated with bright emission and at high power simplifies to $\bar{n}_{b}-\bar{n}_{d}\approx\Lambda\eta$. Here, $\Gamma_{p}$ is the power-dependent polarization rate, $\Lambda=2244\pm 108$ is the cyclicity, $\tau=50\,\mathrm{\mu s}$ is the duration of the readout collection window, and as before $\bar{n}_{b}$ ($\bar{n}_{d}$) are the mean number of counts detected when the qubit is prepared in the spin state associated with bright (dark) emission.
@@
We fit Eq. 6 to the measurement in Fig. 4b. From this fit, we determine efficiency of $\eta=0.992\times 10^{-3}\pm 5.73\times 10^{-6}\approx 0.1\%$; this is lower than we expect from the data in Fig. 3 because of the lower mean count number difference of $\bar{n}_{b}-\bar{n}_{d}\approx 2$ collected photons per cycle at high powers. We attribute this difference to changing conditions of the setup between measurements including drift in alignment and fluctuating polarization of the excitation laser.
@@
Figure 4b shows that the mean number of emitted photons plateaus near a saturation power of $p_{\mathrm{sat}}=313\pm 8\,\mathrm{nW}$, specified at the input to the cryostat. This affirms the straightforward conclusion that readout should be optimized by operating near saturation. Below saturation, emission can be increased by pumping harder. Above saturation, noise may increase with power but signal will not. See Fig. 15a for a measurement of readout fidelity vs power.
@@
Many sources of loss contribute to inefficiency including nonradiative decay, filtering of emission to collect the phonon sideband only, loss between the qubit and detector, and detector inefficiency (see Table 2 for details). However, we estimate the dominant source of loss is the scattering of emission into bulk diamond rather than into the collection path. Although this is the same emitter used in Ref. [30], from fine-tuned optimization using nanopositioners we confirm that this emitter is actually in a mesa structure (see Fig. 11) rather than a photonic nanopillar. Numerical simulations of an SnV- center within this structure predict that only a modest approximately $5\%$ of emitted light is routed to collection (compared to approximately $3\%$ for an emitter in bulk).
@@
This fraction can be greatly increased by use of nanophotonics. For example, at visible wavelengths this number can be increased to approximately 90% using photonic waveguides or photonic crystals [55, 33]. Likewise, inverse-designed grating couplers can extract approximately $25\%$ [56] and tapered fiber coupling up to approximately $90\%$ [57] of photons in a waveguide mode. Since stable, narrow linewidth SnV-’s have already been successfully incorporated into nanophotonic structures [32, 33, 34, 35, 36, 38, 37], utilizing nanophotonics offers a clear path toward an order-of-magnitude higher efficiency, which will advantageously improve readout fidelity at low excitation powers, Fig. 4c.
Improved efficiency also lowers the cyclicity required for single-shot readout (generally $\eta\Lambda>1$), allowing for single-shot readout over expanded magnetic field orientations and strain regimes. For instance, this will enable single-shot readout of highly strained SnV-’s which, in general have lower cyclicity with more sensitivity to magnetic field alignment, but for which microwave spin control can require less drive power [31, 28].
@@
Figure 5: Control of qubit dephasing via weak quantum measurement. (a) A qubit prepared in a superposition state has an initial coherence, specified by the amplitude of its off-diagonal density matrix elements $|\hat{\rho}_{\downarrow\uparrow}|$. This is illustrated on the Bloch sphere for the superposition state (|+|$(\ket{\downarrow}+\ket{\uparrow})/\sqrt{2}$. (b) Measurement causes decoherence at rate $\Gamma_{\phi}$. In a projective measurement the qubit is entirely dephased, but, more generally, after a weak measurement its coherence is reduced to $|\hat{\rho}_{\downarrow\uparrow}^{\prime}|\leq|\hat{\rho}_{\downarrow\uparrow}|$. (c) This interplay is studied by inserting a weak measurement of variable power $p/p_{\mathrm{sat}}$ and variable duration $\tau$ into the center of a dynamical decoupling experiment. (See Fig. 12 for details.) The (projective) readout signal of this experiment measures the ratio $r$ of the remaining coherence to the initial coherence, $r=|\hat{\rho}_{\downarrow\uparrow}^{\prime}|/|\hat{\rho}_{\downarrow\uparrow}|$, which decays exponentially with weak measurement time as $r=\mathrm{exp}\left(-\Gamma_{\phi}\tau\right)$. (d) Selected measurements of $r$ as a function of $\tau$ and $p/p_{\mathrm{sat}}$ and compared to exponential fits. (e) Dephasing rate vs power. Below saturation, the dephasing rate rises linearly with power. Above saturation, it plateaus to $\gamma/4$, the maximum dephasing rate set by the SnV-’s optical decay rate $\gamma$. Data are fit to Eq. 7.